OpenStudy (omgthatgirl):

What is the equation of a circle with center (-7,-3) and radius 2? a. (x+7)^2 + (y+3)^2=4 b. (x-7)^2 + (y-3)^2=4 c. (x-7)^2 + (y-3)^2=2 d. (x+7)^2 +(y+3)^3=2

OpenStudy (igreen):

Equation of a Circle: \(\sf (x - h)^2 + (y - k)^2 = r^2\)

OpenStudy (igreen):

Where \(\sf (h, k)\) is the center.

OpenStudy (igreen):

So in this case: \(\sf h = -7\) \(\sf k = -3\) \(\sf r = 2\)

OpenStudy (igreen):

Can you plug those numbers in? @Omgthatgirl

OpenStudy (omgthatgirl):

yes

OpenStudy (igreen):

Okay, tell me what you get.

OpenStudy (omgthatgirl):

Ok is this how the equation should be (x-7)^2 + (y-3)^2=2^2

OpenStudy (omgthatgirl):

?

OpenStudy (omgthatgirl):

@iGreen